#4) A 30-year-old 200 unit apartment complex is for sale in Canton, GA for $12,000,000. Using the financials below what is the NOI, Cash Flow, Asking Cap Rate, Return on Equity, Debt Coverage Ratio, Loan to Value?

If you want at least a 10% return on your investment how much would you pay for the complex? Is this a good deal for you?

Show your work.

Basic Financial Information:

200 units (#100/2-bedroom/ 1 Bath; #70/1 Bedroom/ 1 bath; #30/ Studio)

Average Rent - $1,100/mt for 2 bedroom; $750/ mt for 1 bedroom; Studio for

$500/mt.

Taxes - $150,000/ yr.

Insurance - $30,000/ yr.                                       General & Admin - $40,000

Amortization – $9,000/ yr.                                  Management – 3% of Effective Gross

Vacancy/ Credit Loss – 10%                                Landscape –$85,000/ yr.

Maintenance – $120,000/ yr.                             Depreciation - $60,000/ yr.

Pest Control - $10,000/ yr.                                  Other Income - $5,000/yr.

Payroll - $250,000/ yr.                                          Utilities - $135,000/ yr.

Capital Improvements – $120,000/ yr.            Advertising - $30,000/ yr.

Equity – $2,400,000

Loan - $9,600,000 @ 8%, Interest only, 7 year balloon

Debt Service - $768,000/ year; $64,000/ month
Answer #1
  1. Calculation of NOI

NOI = (potential rental income – vacancy and credit loss + other income) – operating expenses

Calculation of POTENTIAL RENTAL INCOME

Particular

Units

Rent per month

Number of months

Rent per year

2 bedroom

100

110

12

1320000

1 bedroom

70

750

12

630000

Studio

30

500

12

180000

Calculation of vacancy and credit loss

     Vacancy and Credit loss = 10% of potential rental income

                                                  = 10% of 2130000

                                                    =213000

Calculation of other income

                          Other income = $5000 year

Calculation of operating expenses

   Insurance                                   30000

Maintenance                                 120000

Pest control                                   10000

Payroll                                            250000

General and admin                      40000

Management                                 57660

Landscape                                       85000  

Utilities                                           135000

Taxes (note a)                               30000

Operating expenses                       150000

NOTE

  1. Taxes are assumed as property taxes
  2. Management 3% of effective gross

i.e = 3%(gross rental income+ other income-vacancy and credit loss)

            =3%(2130000+5000-213000)

            =3%(1922000)

              =57660

NOI = 2130000-213000+5000-907660

       = $1014340

(

(c) calculation of cashflow

cashflow from operation

=NOI-capital improvements

=1014340-120000

=894340

Cashflow after financing

=cashflow from operation –financing cost

=894340 – (9600000*8/100)

=894340-768000

=126340

(c) asking cap rate

= net operating income/current market value(sale price of asset)

=1014340/12000000*100

=8.45%

  1. Calculation of NOI

NOI = (potential rental income – vacancy and credit loss + other income) – operating expenses

Calculation of POTENTIAL RENTAL INCOME

Particular

Units

Rent per month

Number of months

Rent per year

2 bedroom

100

110

12

1320000

1 bedroom

70

750

12

630000

Studio

30

500

12

180000

2130000

Calculation of vacancy and credit loss

     Vacancy and Credit loss = 10% of potential rental income

                                                  = 10% of 2130000

                                                    =213000

Calculation of other income

                          Other income = $5000 year

Calculation of operating expenses

   Insurance                                   30000

Maintenance                                 120000

Pest control                                   10000

Payroll                                            250000

General and admin                      40000

Management                                 57660

Landscape                                       85000  

Utilities                                           135000

Taxes (note a)                               30000

Operating expenses                       150000

NOTE

  1. Taxes are assumed as property taxes
  2. Management 3% of effective gross

i.e = 3%(gross rental income+ other income-vacancy and credit loss)

            =3%(2130000+5000-213000)

            =3%(1922000)

              =57660

NOI = 2130000-213000+5000-907660

       = $1014340

(

(c) calculation of cashflow

cashflow from operation

=NOI-capital improvements

=1014340-120000

=894340

Cashflow after financing

=cashflow from operation –financing cost

=894340 – (9600000*8/100)

=894340-768000

=126340

(c) asking cap rate

= net operating income/current market value(sale price of asset)

=1014340/12000000*100

=8.45%

(d) Calculation of Return on Equity

=net income/Shareholder equity

Net income =Rental income +Other income-Operating expenses.

=2130000+5000-907660

=1227340

Shareholder equity is 2400000

ROE=1227340/2400000=.0512:1

Debt coverage ratio

=NOI/ Total debt Services

=1014340/768000

=1.32:1

Loan to value ratio

=Amt of loan/Total value of asset

=9600000/12000000 *100

=80%